Six balls numbered 1, 2, 3, 5, 8, and 13 are placed in a box. A ball is

selected at random, and its number is recorded and then it is replaced. Find

the expected value of the numbers that will occur.


1
Expert's answer
2021-09-02T10:20:53-0400

Since there are 6 numbers with equal probabilities of being selected

P(Selecting any number) "= \frac{1}{6}"

Expectation of the number "= \frac{(1 + 2 + 3 + 5 + 8 + 13)}{6} = \frac{32}{6} = 5.33"